OpenStudy (anonymous):

limx-> 0 (1/x - 1/|x|)

zepdrix (zepdrix):

\[\Large\bf\sf \lim_{x\to0}\frac{1}{x}-\frac{1}{|x|}\]Are we approaching from the left or right of zero? Or that's not specified?

OpenStudy (anonymous):

that's not specified

zepdrix (zepdrix):

\[\Large\bf\sf \frac{1}{|x|}\quad=\quad \frac{1}{x},\qquad x>0\]\[\Large\bf\sf \frac{1}{|x|}\quad=-\frac{1}{x},\qquad x< 0\] So when we're approaching zero from the left side we have,\[\Large\bf\sf \lim_{x\to0^-}\frac{1}{x}-\left(-\frac{1}{x}\right)\]And from the right,\[\Large\bf\sf \lim_{x\to0^+}\frac{1}{x}-\frac{1}{x}\] For the limit to exist (with neither left nor right specificed), the limits from the left and right have to agree.

zepdrix (zepdrix):

Evaluate both of those limits. Determine whether or not you get the same values.

OpenStudy (anonymous):

ok Thank you

zepdrix (zepdrix):

welcome c: